# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

### v3

@@ Abstract
Adaptability to specialised or homogeneous behaviours is critical in cooperative multi-agent reinforcement learning (MARL). Parameter sharing (PS) techniques, common for efficient adaptation, often limit behavioural diversity due to cross-agent gradient interference, which we show can be exacerbated by the coupling of observations and agent IDs. Current remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates. We ask: can shared policies adapt without these complexities? We propose HyperMARL, a PS approach using hypernetworks for dynamic agent-specific parameters, without altering the RL objective or requiring preset diversity levels. HyperMARL’s explicit decoupling of observation- and agent-conditioned gradients empirically reduces policy gradient variance, facilitates shared-policy adaptation (including specialisation), and helps mitigate cross-agent interference. Across diverse MARL benchmarks (up to 20 agents), requiring homogeneous, heterogeneous, or mixed behaviours, HyperMARL achieves competitive performance against key baselines – fully shared, non-parameter sharing, and three diversity-promoting methods – while preserving behavioural diversity comparable to non-parameter sharing. These findings establish HyperMARL as a versatile approach for adaptive MARL. The code is publicly available at https://github.com/KaleabTessera/HyperMARL.
@@ 1 Introduction
Specialist and generalist behaviours are critical to collective intelligence, enhancing performance and adaptability in both natural and artificial systems [47, 38, 39, 22, 45]. In Multi-Agent Reinforcement Learning (MARL), this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [25, 6, 2].
Figure 1: HyperMARL Policy Architecture. Common Agent-ID conditioned shared MARL policy (left) vs HyperMARL (right), which uses a hypernetwork to generate agent-specific policies and decouples observation- and agent-conditioned gradients.
@@
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards [25, 21], role-based allocations [43, 44], specialised architectures [23, 26, 6], or sequential updates [49]. However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates? A key difficulty with FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [10, 49]. We not only empirically validate this hypothesis but also demonstrate a critical insight: this conflict is significantly exacerbated by the common practice of coupling observations with agent IDs within a shared network (Fig. 1 for coupling, Sec. 3.2 for results).
To counteract the challenges posed by this coupling, we propose HyperMARL, a novel parameter sharing method that explicitly decouples observation- and agent-conditioned gradients (Section 4.3) through hypernetworks [18]. Specifically, agent-conditioned hypernetworks, which condition on an agent identifier (or embedding), and generate an agent’s actor and critic weights on the fly (Fig. 1). Hypernetworks have already proved effective for resolving gradient conflicts in multi-task RL [30] and for mitigating catastrophic forgetting in continual learning [42], our work establishes their effectiveness in MARL. We further show that HyperMARL empirically attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Sec. 5.26), suggesting it helps mitigate cross-agent gradient interference in shared architectures.
We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [34], and SMAX [36]—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo [6],HAPPO [49] and Kaleidoscope [26], while achieving NoPS-level behavioural diversity while using a shared architecture.
@@
  • Motivated by evidence that coupling agent IDs with observations can be detrimental, we propose HyperMARL (Sec. 4), a method that leverages agent-conditioned hypernetworks to decouple observation- and agent-conditioned gradients. HyperMARL adaptively learns diverse, homogeneous, or mixed behaviours without altering the RL objective or requiring preset diversity levels.
  • We thoroughly evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents. HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo, HAPPO and Kaleidoscope, while achieving NoPS-level behavioural diversity.
  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6).
@@ 2 Background
In this setting, each agent $i$ receives a partial observation $o_{t}^{i}\in\mathbb{O}^{i}$. These observations are accumulated into an action-observation history $h_{t}^{i}=\left(o_{0}^{i},a_{0}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i}\right)$. Each agent $i$ acts based on their decentralized policies $\pi^{i}(a^{i}|h^{i})$. The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$. The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows, [^1] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity. $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$, where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
@@ 3 Are Independent or Fully Shared Policies Enough?
Standard independent (NoPS) and fully parameter-shared (FuPS) policies face inherent trade-offs in MARL. NoPS allows for uninhibited agent specialisation but can be sample inefficient and computationally expensive. FuPS, often conditioned with an agent ID (FuPS+ID), is more efficient but can struggle when agents must learn diverse behaviours. This section investigates the limitations of these common policy architectures.
@@
Two-Player Payoff matrix
$N$-player Interaction
@@
Two-Player Payoff matrix
$N$-player Interaction
@@
Figure 2: Specialisation and Synchronisation Games. The Specialisation game (left), which encourages distinct actions, and the Synchronisation game (right), where rewards encourage identical actions. Depicted are their two-player payoff matrices (pure Nash equilibria in blue) and $N$-player interaction schematics. While simple in form, these games are challenging MARL benchmarks due to non-stationarity and exponentially scaling observation spaces (temporal version).
To probe these limitations, we introduce two illustrative environments: the Specialisation Game, rewarding distinct actions, and the Synchronisation Game, rewarding identical actions. Both are inspired by prior work [16, 5, 32] and extended here to $N$-agent and temporal settings where agents observe prior joint actions (see Appendix E for full definitions).
@@ 3.1 Limitations of Fully Shared and Independent Policies
Figure 3: Average evaluation reward (mean $\pm$ 95% CI) for temporal Specialisation vs. Synchronisation using REINFORCE (10 seeds). Bold: highest mean, no CI overlap. Neither fully shared nor independent policies consistently achieve the highest mean reward.
@@
Figure 4: Multi-agent policy gradient methods in the Specialisation environment. The FuPS+ID (No State) ablation outperforms FuPS+ID, showing near-orthogonal gradients (4(b)), indicating that observation–ID decoupling is important. HyperMARL (MLP) enables this decoupling while leveraging state information, and achieves better performance and reduced gradient conflict than FuPS+ID.
FuPS without agent IDs provably cannot recover optimal pure Nash equilibria in the non-temporal 2-player Specialisation Game (Proof E.3, App. E). In practice, however, FuPS is often conditioned with agent IDs, and MARL policies must handle complexities beyond static, two-player interactions. We therefore evaluate standard architectures in the temporal $n$-player versions of these games [^2] Results for non-temporal (normal-form) variants are in App. E.5. We compare three standard architectures trained with REINFORCE [46]: 1) NoPS: independent policies ($\pi_{\theta^{i}}(a^{i}|o^{i})$); 2) FuPS: a single shared policy ($\pi_{\theta}(a^{i}|o^{i})$); and 3) FuPS+ID: a shared policy incorporating a one-hot agent ID ($\pi_{\theta}(a^{i}|o^{i},id^{i})$). All use single-layer networks, 10-step episodes, and $10,000$ training steps (further details in App. 6).
@@ 3.2 Why FuPS+ID Fails to Specialise: The Problem of Gradient Conflict
Importance of Observation and ID Decoupling To investigate the effect of entangled observation and ID inputs, we introduce an ablation: FuPS+ID (No State), where the policy $\pi_{\theta}(a^{i}\mid id^{i})$ conditions only on the agent ID, ignoring observations. Surprisingly, FuPS+ID (No State) outperforms standard FuPS+ID in the Specialisation Game for all tested $N$ (Figure 4(a)), even when $N\leq 4$ (where observation spaces are small, suggesting the issue is not merely observation size). Figure 4(b) reveals why: FuPS+ID (No State) shows near-zero gradient conflict (nearly orthogonal gradients), whereas standard FuPS+ID exhibits negative cosine similarities (conflicting gradients).
@@ 4.1 Hypernetworks for MARL
Linear Hypernetworks Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as :
@@ 4.3 Gradient Decoupling in HyperMARL
A core difficulty in FuPS is cross-agent gradient interference [10, 49]. HyperMARL mitigates this by generating each agent’s parameters through a shared hypernetwork, thereby decoupling agent-conditioned and observation-conditioned components of the gradient.
@@
  • Agent-conditioned factor $\mathbf{J}_{i}$. This Jacobian depends only on the fixed embedding $e^{i}$ and the hypernetwork weights $\psi$, therefore, it is deterministic with respect to mini-batch samples (as $e^{i}$ and $psi$ are fixed per gradient step), separating agent identity from trajectory noise.
@@
The crucial structural decoupling in Equation (4) ensures HyperMARL first averages noise per agent (via factor $Z_{i}$) before applying the deterministic agent-conditioned transformation $\mathbf{J}_{i}$. This mitigates gradient interference common in FuPS+ID  [10, 49], where observation noise and agent identity become entangled (see Equation (12) in App. F.2). This is the MARL analogue of the task/state decomposition studied by [37, Eq. 18] in Meta-RL. Section 5.2 empirically verifies the predicted variance drop, and ablations in Section 6 demonstrate that disabling decoupling degrades performance, underscoring its critical role.
@@ 5 Experiments
To address these questions, HyperMARL is evaluated against five representative modern baselines across a carefully selected suite of MARL benchmarks (18 scenarios, up to 20 agents). All experiments use at least 5 seeds (details in App. G.2.1).
@@ 5.1 Experimental Setup
Environments. HyperMARL is evaluated across 18 scenarios from four diverse MARL environments (Dispersion [5], Navigation [5], MAMuJoCo [34], and SMAX [36]) (Table 1). These were specifically chosen to rigorously test performance across varying complexities, agent counts (2 to 20), and distinct behaviours (heterogeneous, homogeneous, or mixed). Full details in Appendix G.1.
Baselines. We evaluate HyperMARL against modern parameter sharing (PS) and diversity-promoting baselines. Core PS comparisons use FuPS+ID and NoPS. For specialisation, we include privileged baselines: DiCo [6] (shared and non-shared weights, preset diversity levels); HAPPO [49] (shared critic, sequential actor updates); and Kaleidoscope [26] (learnable masks, critic ensembles, diversity loss). We use IPPO/MAPPO [12, 48] as the underlying algorithm for all methods except Kaleidoscope (MATD3 backbone due to its intricate off-policy design; see App. H.1 for these results).
@@
Figure 5: Performance and gradient analysis. (a,b) IPPO and MAPPO on Dispersion (20M timesteps) - IQM of Mean Episode Return with 95% bootstrap CIs: Hypernetworks match NoPS performance while FuPS struggle with specialisation. Interval estimates in App. H.2.1. (c) Actor gradient variance: Hypernetworks achieve lower gradient variance than FuPS+ID. (d) Policy diversity (SND with Jensen–Shannon distance): Hypernetworks achieve NoPS-level diversity while sharing parameters.

### v4

@@ Abstract
Adaptive cooperation in multi-agent reinforcement learning (MARL) requires policies to express homogeneous, specialised, or mixed behaviours, yet achieving this adaptivity remains a critical challenge. While parameter sharing (PS) is standard for efficient learning, it notoriously suppresses the behavioural diversity required for specialisation. This failure is largely due to cross-agent gradient interference, a problem we find is surprisingly exacerbated by the common practice of coupling agent IDs with observations. Existing remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates – raising a fundamental question: can shared policies adapt without these intricacies? We propose a solution built on a key insight: an agent-conditioned hypernetwork can generate agent-specific parameters and decouple observation- and agent-conditioned gradients, directly countering the interference from coupling agent IDs with observations. Our resulting method, HyperMARL, avoids the complexities of prior work and empirically reduces policy gradient variance. Across diverse MARL benchmarks (22 scenarios, up to 30 agents), HyperMARL achieves performance competitive with six key baselines while preserving behavioural diversity comparable to non-parameter sharing methods, establishing it as a versatile and principled approach for adaptive MARL. The code is publicly available at https://github.com/KaleabTessera/HyperMARL.
@@ 1 Introduction
Specialist and generalist behaviours are critical to collective intelligence, enhancing performance and adaptability in both natural and artificial systems [52, 43, 44, 23, 50]. In Multi-Agent Reinforcement Learning (MARL) [2], this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [26, 6].
Figure 1: HyperMARL Policy Architecture. Common agent-ID conditioned shared MARL policy (FuPS+ID, left) vs HyperMARL (right), which uses an agent-conditioned hypernetwork to generate agent-specific weights and decouples observation- and agent-conditioned gradients.
@@
This failure of FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [11, 54]. We not only empirically validate this hypothesis but also demonstrate a critical insight: this conflict is significantly exacerbated by the common practice of coupling observations with agent IDs within a shared network (Fig. 1 for coupling, Sec. 3.2 for results).
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards [26, 22], role-based allocations [48, 49], specialised architectures [24, 28, 6], sequential updates [54], and sharing parameters within clusters of agents [11]. However, these remedies introduce their own intricacies: they often alter the learning objective, require prior knowledge of optimal diversity levels, necessitate maintaining agent-specific parameters or require sequential updates. This raises a fundamental question: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates?
Guided by our observation-ID coupling insight, we propose HyperMARL, a novel agent-conditioned hypernetwork [19] architecture. HyperMARL generates per-agent weights on the fly (Fig. 1) and explicitly decouples observation- and agent-conditioned gradients (Sec. 4.3). This choice is motivated by hypernetworks’ proven effectiveness at resolving gradient conflicts in multi-task RL [33] and continual learning [47]. Our work establishes their effectiveness for the problem of cross-agent interference in MARL. Indeed, HyperMARL empirically attains lower policy gradient variance than FuPS, and we show this decoupling is critical for specialisation (Sec. 5.26.1), confirming its role in mitigating interference.
We validate HyperMARL on diverse MARL benchmarks – including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [37], SMAX [41], and Blind-Particle Spread (BPS) [11] – across environments with two to thirty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as Diversity Control (DiCo) [6], Heterogeneous-Agent Proximal Policy Optimisation (HAPPO) [54], Kaleidoscope [28] and Selective Parameter Sharing (SePS) [11], while achieving NoPS-level behavioural diversity while using a shared architecture.
@@
  • We identify that cross-agent gradient interference in shared policies is critically exacerbated by the common practice of coupling agent IDs with observations (Sec. 3.2).
  • We propose HyperMARL (Sec. 4), an agent-conditioned hypernetwork architecture, to test the hypothesis that explicitly decoupling these gradients enables adaptive (diverse, homogeneous, or mixed) behaviours without the complexities of prior remedies (e.g., altered objectives, preset diversity levels, or sequential updates).
  • Our extensive evaluation (Sec. 5) across 22 diverse scenarios (up to 30 agents) shows HyperMARL achieves competitive returns against six strong baselines, while achieving NoPS-level behavioural diversity. We further show this decoupling is empirically linked to reduced policy gradient variance and is critical for specialisation (Sec. 5.2; Sec. 6.1).
@@ 2 Background
In this setting, each agent $i$ receives a partial observation $o_{t}^{i}\in\mathbb{O}^{i}$. These observations are accumulated into an action-observation history $h_{t}^{i}=\left(o_{0}^{i},a_{0}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i}\right)$. Each agent $i$ acts based on their decentralised policies $\pi^{i}(a^{i}|h^{i})$. The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$. The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows, [^1] We use simplified notation here omitting explicit dependence on state transitions and distributions for brevity. $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$, where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
@@ 3 Are Independent or Fully Shared Policies Enough?
Standard independent (NoPS) and fully parameter-shared (FuPS) policies face inherent trade-offs in MARL. NoPS allows for uninhibited agent specialisation but can be sample inefficient and computationally expensive. FuPS, often conditioned with an agent ID (FuPS+ID), is more efficient but can struggle when agents must learn diverse behaviours [11]. This section investigates the limitations of these common policy architectures.
@@
Two‐Player Payoff matrix
NN‐player Interaction
@@
Two‐Player Payoff matrix
NN‐player Interaction
@@
Figure 2: Specialisation and Synchronisation Games. The Specialisation game (left), which encourages distinct actions, and the Synchronisation game (right), where rewards encourage identical actions. Depicted are their two-player payoff matrices (pure Nash equilibria in blue) and NN-player interaction schematics. While simple in form, these games are challenging MARL benchmarks due to non-stationarity and exponentially scaling observation spaces (temporal version).
To probe these limitations, we introduce two illustrative environments: the Specialisation Game, rewarding distinct actions, and the Synchronisation Game, rewarding identical actions. Both are inspired by prior work [10, 17, 5, 35] and extended here to $N$-agent and temporal settings where agents observe prior joint actions (see Appendix E for full definitions).
@@ 3.1 Limitations of Fully Shared and Independent Policies
Table 1: Average evaluation reward (mean $\pm$ 95% CI) for temporal Specialisation vs. Synchronisation using REINFORCE (10 seeds). Bold: highest mean, no CI overlap. Neither fully shared nor independent policies consistently achieve the highest mean reward.
@@
Figure 3: Multi-agent policy gradient methods in the Specialisation environment. The FuPS+ID (No State) ablation outperforms FuPS+ID, showing near-orthogonal gradients (3(b)), indicating that observation–ID decoupling is important. HyperMARL (MLP) enables this decoupling while leveraging state information, and achieves better performance and reduced gradient conflict than FuPS+ID.
FuPS without agent IDs provably cannot recover optimal pure Nash equilibria in the non-temporal 2-player Specialisation Game (Proof E.3, App. E). In practice, however, FuPS is often conditioned with agent IDs, and MARL policies must handle complexities beyond static, two-player interactions. We therefore evaluate standard architectures in the temporal $n$-player versions of these games [^2] Results for non-temporal (normal-form) variants are in App. E.5. We compare three standard architectures trained with REINFORCE [51]: 1) NoPS: independent policies ($\pi_{\theta^{i}}(a^{i}|o^{i})$); 2) FuPS: a single shared policy ($\pi_{\theta}(a^{i}|o^{i})$); and 3) FuPS+ID: a shared policy incorporating a one-hot agent ID ($\pi_{\theta}(a^{i}|o^{i},id^{i})$). All use single-layer networks, 10-step episodes and $10,000$ training steps (further details in Table 9 in App. I).
@@ 3.2 Why FuPS+ID Fails to Specialise: The Problem of Gradient Conflict
Importance of Observation and ID Decoupling. To investigate the effect of entangled observation and ID inputs, we introduce an ablation: FuPS+ID (No State), where the policy $\pi_{\theta}(a^{i}\mid id^{i})$ conditions only on the agent ID, ignoring observations. Surprisingly, FuPS+ID (No State) outperforms standard FuPS+ID in the Specialisation Game for all tested $N$ (Figure 3(a)), even when $N\leq 4$ (where observation spaces are small, suggesting the issue is not merely observation size). Figure 3(b) reveals why: FuPS+ID (No State) shows near-zero gradient conflict (nearly orthogonal gradients), whereas standard FuPS+ID exhibits negative cosine similarities (conflicting gradients).
@@ 4.1 Hypernetworks for MARL
Linear Hypernetworks Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as follows [^3] For conciseness we only show the policy parameters in this section.:
@@ 4.3 Gradient Decoupling in HyperMARL
A core difficulty in FuPS is cross-agent gradient interference [11, 54]. HyperMARL mitigates this by generating each agent’s parameters through a shared hypernetwork, thereby decoupling agent-conditioned and observation‑conditioned components of the gradient.
@@
  • Agent-conditioned factor $\mathbf{J}_{i}$. This Jacobian depends only on the fixed embedding $e^{i}$ and the hypernetwork weights $\psi$, therefore, it is deterministic with respect to mini-batch samples (as $e^{i}$ and $\psi$ are fixed per gradient step), separating agent identity from trajectory noise.
@@
The crucial structural decoupling in Equation (4) ensures HyperMARL first averages noise per agent (via factor $Z_{i}$) before applying the deterministic agent-conditioned transformation $\mathbf{J}_{i}$. This mitigates gradient interference common in FuPS+ID  [11, 54], where observation noise and agent identity become entangled (see Equation (12) in App. F.2). This is the MARL analogue of the task/state decomposition studied by [42, Eq. 18] in Meta-RL. Section 5.2 empirically verifies the predicted variance drop, and ablations in Section 6.1 demonstrate that disabling decoupling degrades performance, underscoring its critical role.
@@ 5 Experiments
To address these questions, HyperMARL is evaluated against six representative modern baselines across a carefully selected suite of MARL benchmarks (22 scenarios, up to 30 agents). All experiments use at least 5 seeds (details in App. G.2.1).
@@ 5.1 Experimental Setup
BPS 15–30 Discrete Hetero.
Environments. HyperMARL is evaluated across 22 scenarios from five diverse MARL environments (Dispersion [5], Navigation [5], MAMuJoCo [37], SMAX [41], BPS [11]) (Table 2). These were specifically chosen to rigorously test performance across varying complexities, agent counts (2 to 30), and distinct behaviours (heterogeneous, homogeneous, or mixed). Full details in Appendix G.1.
Baselines. We evaluate HyperMARL against modern parameter sharing (PS) and diversity-promoting baselines. Core PS comparisons use FuPS+ID and NoPS. For specialisation, we include privileged baselines: DiCo [6] (shared and non-shared weights, preset diversity levels); HAPPO [54] (shared critic, sequential actor updates); Kaleidoscope [28] (learnable masks, critic ensembles, diversity loss); and SePS (pre-training phase, agent clustering). We use IPPO/MAPPO [13, 53] as the underlying algorithm for all methods except Kaleidoscope and SePS (see App. H.1 and App. H.2 for these results).
@@
Figure 4: Performance and gradient analysis. (a,b) IPPO and MAPPO on Dispersion (20M timesteps) - IQM of Mean Episode Return with 95% bootstrap CIs: Hypernetworks match NoPS performance while FuPS struggle with specialisation. Interval estimates in App. H.3.1. (c) Actor gradient variance: Hypernetworks achieve lower gradient variance than FuPS+ID. (d) Policy diversity (SND with Jensen–Shannon distance): Hypernetworks achieve NoPS-level diversity while sharing parameters.